# Strict area law entanglement versus chirality

### v1

@@ C.2.2 Properties of $\Sigma(A,B,C)$
We first study the deformations that are away from the triple intersection points, shown in Fig. 3 (Left). The derivation utilizes Markov conditions $I(x:B|A)=0,I(y:A|B)=0,I(z:A|B)=0$ from Fact 2.1999The derivation of the last Markov condition from Fact 2.1 is explained later..
@@
With a $\textsf{U}(1)$ symmetric state $\left|\Psi\right\rangle$ that satisfies bulk A1 in the setup described in the main text, on a tripartite region $ABC$ shown in Fig. 1 (Right), we shall consider the quantity
$\langle Q_{BC}^{2}\rangle(t)\equiv\left\langle\Psi(t)\right|Q_{BC}^{2}\left|\Psi(t)\right\rangle,$ (E.3)
where $\left|\Psi(t)\right\rangle=\mathcal{I}_{AB}(t)\left|\Psi\right\rangle$. We shall show that $\langle Q_{BC}^{2}\rangle(t)$ as a function of $t$ is unbounded in both directions. Since $Q_{BC}^{2}$ is a positive operator, this implies $\langle Q_{BC}^{2}\rangle(t)$ must be infinite, i.e., $Q_{BC}$ has eigenvalue(s) being infinite. Since $BC$ is a set of a finite number of sites, the result that $Q_{BC}=\sum_{v\in BC}Q_{v}$ has infinite eigenvalue(s) contradicts the setup, i.e. the on-site $\textsf{U}(1)$ symmetry is not well-defined.
Let’s compute
@@
$\langle Q_{BC}^{2}\rangle(t)=\langle Q_{BC}^{2}\rangle+2\sigma_{xy}t,\quad t\in\mathbb{R}.$ (E.6)
If $\sigma_{xy}\neq 0$, this result implies that $\langle Q_{BC}^{2}\rangle(t)$ is unbounded from both sides as a function of $t$.

### v2

@@ C.2.2 Properties of $\Sigma(A,B,C)$
We first study the deformations that are away from the triple intersection points, shown in Fig. 3 (Left). The derivation utilizes Markov conditions $I(x:B|A)=0,I(y:A|B)=0,I(z:A|B)=0$ from Fact 2.1131313The derivation of the last Markov condition from Fact 2.1 is explained later..
@@
Consider a $\textsf{U}(1)$ symmetric state $\left|\Psi\right\rangle$ on a two-dimensional lattice in a tensor product Hilbert space with finite local dimension as described in the main text. (1) On one hand, because the local Hilbert space $\mathcal{H}_{X}$ of any region $X$ with a finite number of sites is finite dimensional, the local charge operator $Q_{X}$ should have a bounded spectrum and $\langle Q_{X}^{2}\rangle$ should be finite. (2) On the other hand, with the conditions that $\left|\Psi\right\rangle$ satisfies bulk A1 and have non-zero $\sigma_{xy}$ from $\Sigma(A,B,C)$ on a tripartite region $ABC$ shown in Fig. 1 (Right), we can show that
$\langle Q_{BC}^{2}\rangle(t)$ $\equiv\left\langle\Psi(t)\right|Q_{BC}^{2}\left|\Psi(t)\right\rangle$ (E.3)
$=\langle Q_{BC}^{2}\rangle+2\sigma_{xy}t,\quad t\in\mathbb{R}.$ (E.4)
where $\left|\Psi(t)\right\rangle=\mathcal{I}_{AB}(t)\left|\Psi\right\rangle$. This implies $\langle Q_{BC}^{2}\rangle(t)$ as a function of $t$ is unbounded in both directions. Since $Q_{BC}^{2}$ is a positive operator and $\langle Q_{BC}^{2}\rangle(t)\geq 0$, we can conclude $\langle Q_{BC}^{2}\rangle$ must be infinite and hence the spectrum of $Q_{BC}$ is unbounded, where $BC$ contains only finite number of sites. Therefore, we have obtained a contradiction.
@@
and obtain Eq. (E.4).
@@
Appendix F Generalizations of the no-go theorems
The no-go theorems 3.1 and 3.2 both involve the finite local Hilbert space dimensions condition. This condition can be relaxed into the condition that $\left|\Psi\right\rangle$ has finite local entanglement entropies and finite local charge fluctuations (i.e. $(S_{A})_{\left|\Psi\right\rangle}<\infty$ and $\langle\Psi|Q_{A}^{2}|\Psi\rangle-\langle\Psi|Q_{A}|\Psi\rangle^{2}<\infty$ for any local region $A$). [^14] We left it a question on if the finiteness of entropy implies the finiteness of charge fluctuation.
Preliminaries: We shall consider a state $\left|\Psi\right\rangle$ with well-defined local reduced density matrices $\rho_{A}$ and modular Hamiltonians $K_{A}=-\ln\rho_{A}$, and their von Neumann entropies are finite $S(\rho_{A})<\infty$, where $A$ stands for any regions with finite number of lattice sites. First of all, one can straightforwardly define instantaneous modular flows in the same manner as in the main text. We shall assume that the modular flow gives physical quantities as a smooth function of $t$. Moreover, the implications of the bulk A1 condition still holds, which we explicitly explain in the following:
  • Fact 2.1 still holds, since its proof [13] only uses strong subadditivity of entanglement entropies, which is true regardless of the local Hilbert space dimension [39].
  • Lemma 4.2 is still valid and therefore, one can still do Markov decomposition move. This is because if $\left|\Psi\right\rangle$ has finite local entanglement entropies, even if it is in an infinite dimensional Hilbert space, the structure theorem Eq. (B.2) of a Markov state $\rho_{ABC}$ from $\left|\Psi\right\rangle$ still holds [Jen2006], and proof of Lemma 4.2 given in App. B is still applicable.
  • One can still decompose the modular Hamiltonian of a Markov state, namely if $I(A:C|B)_{\left|\Psi\right\rangle}=0$, then
    $K_{ABC}\left|\Psi\right\rangle=(K_{AB}+K_{BC}-K_{B})\left|\Psi\right\rangle,$ (F.1)
    where $K_{\bullet}$ are modular Hamiltonians of $\rho_{\bullet}$ from $\left|\Psi\right\rangle$. This equation can be obtained by taking the first derivative with respect to $t$ at $t=0$ for the Markov flow equation $\mathcal{I}_{ABC}(t)\left|\Psi\right\rangle=\mathcal{I}_{AB}(t)\mathcal{I}_{BC}(t)\mathcal{I}_{B}(-t)\left|\Psi\right\rangle,\forall t\in\mathbb{R}$.
In summary, if $\left|\Psi\right\rangle$ satisfies bulk A1 and has finite local entanglement entropies, one can still do Markov decomposition move and decompose the modular Hamiltonian as in Eq. (F.1) for a Markov state.
Now we discuss the modified no-go theorems:
Theorem F.1.
There is no quantum many-body state $\left|\Psi\right\rangle$ in a tensor product Hilbert space with finite local entanglement entropies satisfying bulk A1 with non-zero $c_{-}$ from $J(A,B,C)_{\left|\Psi\right\rangle}$ using Eq. (2.4).
Theorem F.2.
There is no quantum many-body state $\left|\Psi\right\rangle$ with on-site $\textsf{U}(1)$ symmetry in a tensor product Hilbert space with finite local entanglement entropies and finite local charge fluctuations satisfying bulk A1 with non-zero $\sigma_{xy}$ from $\Sigma(A,B,C)_{\left|\Psi\right\rangle}$ using Eq. (2.5).
The proofs of Theorem F.1 and Theorem F.2 are vastly overlapped with the proofs of Theorem 3.1 and Theorem 3.2. Below we explain how the original proofs can be applied in this context:
Proof sketch of Theorem F.1: The proof is still by contradiction. Let $\left|\Psi\right\rangle$ be a state with all the properties in Theorem F.1.
(1) Firstly, Lemma 5.1 is still valid and the proof is the same as the one in App. D: The step 1 in the proof of Lemma 5.1 can be repeated because one can still apply Markov decomposition move, which is the only non-trivial ingredient. For step 2, the main ingredients are strong subadditivity and decomposition of modular Hamiltonian of a Markov state Eq. (F.1), which are also applicable as we discussed above, therefore the step 2 in the proof of Lemma 5.1 is also applicable.
(2) Once Lemma 5.1 is proved, then one can repeat the same derivation mentioned in the main body of the proof of Theorem 3.1 in the main text, and conclude
$(S_{BC})_{\left|\Psi(t)\right\rangle}=(S_{BC})_{\left|\Psi\right\rangle}+{\pi c_{-}\over 3}t,\quad\forall t\in\mathbb{R}.$ (F.2)
Since $(S_{BC})_{\left|\Psi(t)\right\rangle}\geq 0,\forall t\in\mathbb{R}$ and one can take $\pi c_{-}t/3\to-\infty$ by letting $t\to-\infty$ (or +$\infty$+ ∞) if $c_{-}>0$ (or $<0$), $(S_{BC})_{\left|\Psi\right\rangle}$ must be infinite, which contradicts to the finite local entropies assumption of $\left|\Psi\right\rangle$.
Proof sketch of Theorem F.2: The proof in App. E can be repeated for the same reasons given above, as it uses the same ingredients (Markov decomposition move and Eq. (F.1) for Markov state). Similar to the proof above, once one repeats the derivations and obtains
$\Sigma(A,B,C)_{\left|\Psi(t)\right\rangle}=\sigma_{xy},\quad\forall t\in\mathbb{R},$ (F.3)
where $\left|\Psi(t)\right\rangle=\mathcal{I}_{AB}(t)\left|\Psi\right\rangle$. Moreover, one can show that the charge fluctuation on $\left|\Psi(t)\right\rangle$ on $BC$, defined as
$\sigma_{BC}(t)\equiv\langle\Psi(t)|Q_{BC}^{2}|\Psi(t)\rangle-\langle\Psi(t)|Q_{BC}|\Psi(t)\rangle^{2}$ (F.4)
satisfies
${d\sigma_{BC}(t)\over dt}=2\Sigma(A,B,C)_{\left|\Psi(t)\right\rangle}$ (F.5)
$+$ $2{\bf i}\langle\Psi(t)|Q_{BC}|\Psi(t)\rangle\cdot\langle\Psi(t)|[K_{AB}(t),Q_{BC}]|\Psi(t)\rangle$
$=$ $2\Sigma(A,B,C)_{\left|\Psi(t)\right\rangle},$
where $K_{AB}(t)$ is the modular Hamitonian from $\left|\Psi(t)\right\rangle$ (which is equal to $K_{AB}$) and the term in the second line vanishes because
$\begin{split}&\langle\Psi(t)|[K_{AB}(t),Q_{BC}]|\Psi(t)\rangle\\ =&\langle\Psi(t)|[K_{AB}(t),Q_{B}]|\Psi(t)\rangle+\langle\Psi(t)|[K_{AB}(t),Q_{C}]|\Psi(t)\rangle\\ =&\langle\Psi(t)|[K_{\overline{AB}}(t),Q_{B}]|\Psi(t)\rangle\\ =&0.\end{split}$ (F.6)
Therefore, we obtain
$\sigma_{BC}(t)=\sigma_{BC}+2\sigma_{xy}t,\quad\forall t\in\mathbb{R},$ (F.7)
where $\sigma_{BC}$ is the charge fluctuation computed on $\left|\Psi\right\rangle$. Since $\sigma_{BC}(t)\geq 0$ and one can take $2\sigma_{xy}t\to-\infty$ if $\sigma_{xy}\neq 0$, then $\sigma_{BC}$ must be infinite, which contradicts to the finite charge fluctuations condition.